A patient is given an injection with a hypodermic needle 3.2 cm long and 0.26 mm in diameter.

Part A

Assuming the solution being injected has the same density and viscosity as water at 20 ∘C, find the pressure difference needed to inject the solution at the rate of 2.0 g/s .
Answer #1

Length of needle L=3.2 cm = 0,032 m

Radius of needle r = 0.13 mm = 0.13 * 10-3 m

Viscosity of solution n = 0.00101 Pa.s

Density of solution p= 1000 kg/m3 = 10g/m

Flow rate of solution is Q=AV Am/At 2.0g/s 106 g/m3 03 = 2.0 * 10-6 m3/s 106

Flow rate Q = TAPA Q= 8nL

Pressure difference 8LQ IP = 74.

ΔΡ - 8 % 0.00101 % 0.032 * 2.0 * 10-6 π(0.13 και 10-34 -= 5.76 και 105 Pa